OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points R(3, 3), S(-6, -6) _ _ _ = _ _ _ _ _ _

OpenStudy (anonymous):

First you have to get the slope. (Y2-Y1)/(X2-X1) In this case it would be 1. Then use slope point form. y-y1=m(x-x1) So you would get y -3 = 1(x-3). which is equal to y=x. So standard form would be -x+y=0. I think that is it.

OpenStudy (anonymous):

thank you can you help me with one more please

OpenStudy (anonymous):

Yeah just write it.

OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):

Natasha I will give you the steps to do.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

First thing you do is use slope formula and get the slope between the two points. After you get the slope you plug it into slope point form. after you plug it into slope point form, you solve for y. One solved for y, you move the x value to the other side (with y) And voilla you're done.

OpenStudy (anonymous):

x-y=4 ?

OpenStudy (anonymous):

Close, it is supposed to be x+y=0 I think.

OpenStudy (anonymous):

thank you